OpenStudy (anonymous):

\lim_{x \rightarrow \infty} \left( \sqrt{x^2 + 9 x + 1} -x \right) =

OpenStudy (anonymous):

\lim_{x \rightarrow \infty} \left( \sqrt{x^2 + 9 x + 1} -x \right)